OpenStudy (anonymous):

CAN ANY ONE there is a groove which is in a cylindrical form and is made in a cylindrical piece of wood of length 17.5cm. The groove is filled with graphite and the cylindrical graphite block is of same length as that of wooden piece. If the diameter of the graphite cylinder is 3mm and outer diameter of the wooden cylinder is 10mm. find the volume of the wood and volume of the graphite used

OpenStudy (amistre64):

theres not enough information that i can see

OpenStudy (amistre64):

there is a groove which is in a cylindrical form and is made in a cylindrical piece of wood of length 17.5cm. is the groove shaped like a cylindar? how can you make a cylindar shaped groove? do they mean a spiral groove around a core, like a screw? The groove is filled with graphite and the cylindrical graphite block is of same length as that of wooden piece. If the diameter of the graphite cylinder is 3mm and outer diameter of the wooden cylinder is 10mm. find the volume of the wood and volume of the graphite used if got no idea what this is trying to picture

OpenStudy (kropot72):

The volume of the wooden cylinder without the groove is \[\frac{\pi 10^{2}}{4}\times 175\ mm ^{3}\] The vloume of the groove is \[\frac{\pi 3^{2}}{4}\times\ 175\ mm ^{3}\] Volume of wood will be the difference of the two values of volume.

OpenStudy (kropot72):

volume*

OpenStudy (anonymous):

volume of graphite

OpenStudy (anonymous):

r u sure that volume of groove u have calculated is not volume of graphite

OpenStudy (anonymous):

diamter of graphite is given

OpenStudy (kropot72):

The volume of the graphite is the same as the volume of the groove. The question states that the groove is "filled with graphite". Volume of groove = volume of graphite = \[\frac{\pi \times 3^{2}}{4}\times 175\ mm ^{3}\]